OpenStudy (anonymous):

(3p+9)^2 9p^2+54p+81

OpenStudy (anonymous):

\[(3p+9)^{2}9p ^{2}+54p+81\]

OpenStudy (anonymous):

27(3p^4+18p^3+27p^2+2p+3)

OpenStudy (anonymous):

That can't be the answer

OpenStudy (anonymous):

The answer is: 81p^4+486p^3+724p^2+54p+81 but how??

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

oh nvm I made a mistake

OpenStudy (anonymous):

Yeah my teacher did the problem. I'm trying to understand the problem of how she got the answer in order for me to do the rest of these problems that are basically the same like this

OpenStudy (anonymous):

(3 p+9)^2 9 p^2+54 p+81 expand (3 p+9)^2 w/FOIL (3 p+9) (3 p+9) = (3 p) (3 p) + (3 p) (9) + (9) (3 p) + (9) (9) = 9 p^2+27 p+27 p+81 = 81+54 p+9 p^2: 81+54 p+9 p^2 9 p^2+54 p+81 distribute 9 p^2 over 9 p^2+54 p+81 9 p^2 (9 p^2+54 p+81) = 9 p^2×81+9 p^2×54 p+9 p^2×9 p^2: 9 p^2×81+9 p^2×54 p+9 p^2×9 p^2+54 p+81 multiply 9 and 81 together 9×81 = 729: 729 p^2+9 p^2×54 p+9 p^2×9 p^2+54 p+81 combine products of like terms 9 p^2×54 p = 9 p^(2+1)×54: 729 p^2+9 p^(2+1)×54+9 p^2×9 p^2+54 p+81 evaluate 2+1 2+1 = 3: 729 p^2+9 p^3×54+9 p^2×9 p^2+54 p+81 multiply 9 and 54 together 9X54 = 486: 729 p^2+486 p^3+9 p^2×9 p^2+54 p+81 ccombine products of like terms 9 p^2×9 p^2 = 9×9 p^4: 729 p^2+486 p^3+9×9 p^4+54 p+81 muultiply 9 and 9 together 9×9 = 81: 729 p^2+486 p^3+81 p^4+54 p+81